OpenStudy (anonymous):

Let f(x)=(8x^3-1)^3 and h(x)=x^3. Given that f(x)=(h of g)(x), find g(x)

OpenStudy (anonymous):

(8x^3-1)^3 =g(x)^3 g(x)=8x^3-1